OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –3(x – 10)2 + 5? Domain: All Real Numbers Range: y 5 Domain: All Real Numbers Range: y 5 Domain: All Real Numbers Range: y 10 Domain: All Real Numbers Range: y 10 I think it's the second one

OpenStudy (anonymous):

\[\ge \le \ge \le \] there should be each of these in between y and the number in this order ^^

hero (hero):

Did you just guess an answer or did you actually calculate a series of steps to get a possible result?

OpenStudy (anonymous):

i chose the answer that i thought was correct

OpenStudy (anonymous):

Good

hero (hero):

You can't just choose what you think is correct.

hero (hero):

You need to use logic and algebra to get the result. Anything otherwise is just guessing.

OpenStudy (anonymous):

thats obviously why im asking for help. @MAXXECUTOR do you think it's right?

OpenStudy (anonymous):

You Are Always Right @LaurenAshley1201

OpenStudy (anonymous):

@MAXXECUTOR haha i'm serious

hero (hero):

You offered a possible result. So I am trying to see if you just need help verifying the answer if you need help with the steps. But my main interest is ensuring you know the steps.

OpenStudy (anonymous):

verifying

OpenStudy (anonymous):

closing this question

hero (hero):

Yeah, your guess is correct because you still haven't explained how you got it.